# Learning from Imperfect Demonstrations with Self-Supervision for Robotic Manipulation

### v2

@@ I Introduction
Our key insight is motivated by the observation: Many steps in a failed trajectory are of high quality despite the overall task failure. For example, when attempting to open a cabinet, the robotic arm may successfully grasp the handle but fail to open the cabinet fully. In such cases, the failed data leading up to the successful grasp is valuable and should be regarded as successful data. More importantly, the expert data from task success certainly contains data that succeeded in grasping the handler, so these two pieces of data will be very similar, providing hints as to how to select high-quality imperfect data.
To this end, we propose a novel Self-Supervised Data Filtering framework (SSDF) for robotic manipulation, which extends the training dataset using high-quality imperfect data filtered based on the similarity between the failed trajectory segment and the expert dataset. Specifically, SSDF contains three main steps. In step 1, we pre-train a multi-modality transformer network using the mixed demonstrations in a self-supervised manner to extract representative features. Inspired by pertaining process in BERT [28], we propose three self-supervised objectives to enhance the pre-trained transformer, including 1) Masked Transition Prediction, 2) Transition Reconstruction, and 3) Action Autoregression. These approaches enable the transformer to take on different roles like a forward dynamic model, inverse dynamic model, and behavior cloning model and thus enhance its ability to extract effective features. In step 2, we aim to identify the useful imperfect demonstrations according to the quality scores and use them to augment the original expert dataset. Thus, we leverage the pre-trained transformer in step 1 to extract the features and then generate the quality scores by calculating the similarity between the extracted features of the imperfect and expert demonstrations for each state-action pair in the imperfect demonstrations. In step 3, we perform weighted behavior cloning for the high-quality imperfect data, and the original expert dataset can be used for any IL method. To differentiate the quality of imperfect demonstrations and ensure their beneficial contribution to the downstream policy, we propose the quality scores calculated in step 2 as weights for the behavior cloning loss. Our contributions are as follows.
@@ II Related Work
Imitation Learning has enormous potential to meet critical safety and cost desiderata for real-world applications. As an important class of methods, Behavior Cloning (BC) [14] can be used in the offline setting seamlessly and gain comparable performance compared to GAIL methods [29]. IBC [30], ACT [31], and Diffusion Policy [8] further enhance BC using the network with better expressive capability and advanced training objectives. However, most of these works are based on the assumption that all demonstrations are optimal, thus dropping their performance with imperfect demonstrations.

### v3

@@ I Introduction
Our key insight is motivated by the observation: Many steps in a failed trajectory are of high quality despite the overall task failure. For example, when attempting to open a cabinet, the robotic arm may successfully grasp the handle but fail to open the cabinet. In such cases, the failed data leading up to the successful grasp is still valuable. More importantly, the expert data from task success certainly contains data that succeeded in grasping the handler, so these two pieces of data will be very similar, providing hints as to how to select high-quality imperfect data.
To this end, we propose a novel Self-Supervised Data Filtering framework (SSDF) for robotic manipulation, which extends the training dataset using high-quality imperfect data filtered based on the similarity between the failed trajectory segment and the expert dataset. Specifically, SSDF contains three main steps. In step 1, we pre-train a multi-modality transformer network using the mixed demonstrations in a self-supervised manner to extract representative features. Inspired by pertaining process in BERT [28], we propose three self-supervised objectives to enhance the pre-trained transformer, including 1) Masked Transition Prediction, 2) Transition Reconstruction, and 3) Action Autoregression. These approaches enable the transformer to take on different roles like a forward dynamic model, inverse dynamic model, and behavior cloning model and thus enhance its ability to extract effective features. In step 2, we aim to identify the useful imperfect demonstrations according to the quality scores and use them to augment the original expert dataset. Thus, we leverage the pre-trained transformer in step 1 to extract the features and then generate the quality scores by calculating the similarity between the extracted features of the imperfect and expert demonstrations for each state-action pair in the imperfect demonstrations. In step 3, we perform weighted behavior cloning for the high-quality imperfect data, and the original expert dataset can be used for any IL method. To differentiate the quality of imperfect demonstrations and ensure their beneficial contribution, we propose the quality scores calculated in step 2 as weights for the behavior cloning loss. Our contributions are as follows.
@@ II Related Work
Imitation Learning has enormous potential to meet critical safety and cost desiderata for real-world applications. As an important class of methods, Behavior Cloning (BC) [14] can be used in the offline setting seamlessly and gain comparable performance compared to GAIL methods [29]. IBC [30], ACT [31], and Diffusion Policy [8] further enhance BC using the network with better expressive capability and advanced training objectives. However, most works are based on the assumption that all demonstrations are optimal, thus dropping their performance with imperfect data. In addition, when scaling up the size of datasets in the real world, such as OpenX [32]and RoboMIND [33], it becomes harder to guarantee that the datasets are all optimal as they get larger.